OpenStudy (xguardians):

Evaluate the function. Fine p(1/2) given p(x) = -13x^3 + 5x^2 - 3x + 2 Award is medal, my last question for the night :)

jimthompson5910 (jim_thompson5910):

everywhere you see an x, replace it with 1/2

jimthompson5910 (jim_thompson5910):

then evaluate/simplify

jimthompson5910 (jim_thompson5910):

p(x) = -13x^3 + 5x^2 - 3x + 2 p(1/2) = -13(1/2)^3 + 5(1/2)^2 - 3(1/2) + 2 p(1/2) = ???

OpenStudy (xguardians):

So. -13(1/2)^3?

jimthompson5910 (jim_thompson5910):

don't forget the parenthesis

jimthompson5910 (jim_thompson5910):

that's part of it

OpenStudy (anonymous):

yeah, x is the input so you replace 1/2 anywhere you see an x

OpenStudy (anonymous):

and then solve

OpenStudy (anonymous):

-13(1/2)^3+5(1/2)^2-3(1/2)+2

OpenStudy (xguardians):